OpenStudy (rheclife304):

What is the distance between points (2, 2) and (1, 1)? 1.32 1.57 1.41 none of the above

OpenStudy (welshfella):

do you know tstance formula for 2 points given thgeire coordinates?

OpenStudy (welshfella):

* the distance formula

OpenStudy (rheclife304):

uhm no maybe because Kanawha County hires pornstars to be math teachers and i got kicked out of class all 9th grade year

OpenStudy (anonymous):

\[\sqrt{1^2+1^2}=\sqrt{2}=1.41 \]

OpenStudy (rheclife304):

thank you for actually helping me

OpenStudy (rheclife304):

What is the distance between points (-5, 9) and (3, -3)? 15 15.6 14.42 12. 5

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (rheclife304):

your welcome idk this stuff my school do care if you fail they just move you up instead of letting you pass

OpenStudy (anonymous):

\[\sqrt{(-3-9)^2+(3--5)^2}=\sqrt{208}=4 \sqrt{13}=14.422 \]